OpenStudy (anonymous):

I really need help with finding the average rate of change. Can someone please help me??

OpenStudy (anonymous):

post the problem and I shall try!

OpenStudy (mathmale):

Sure, Jon. What's the function? and on what interval are we trying to find the average rate of change?

OpenStudy (anonymous):

Use the graph below to answer the following question. What is the average rate of change from x = -1 to x = 1? A. -4 B. -2 C. 2 D. 4 I know the average rate of change formula is --> \[A(x)=\frac{ f(x)-f(a) }{ x-a }\] But I have no clue how to plug in the numbers or the process.

OpenStudy (mathmale):

Jon: The interval of interest is [-1,1]. Please evaluate f(x) at each value. Yes, I know that the function is not given as an algebraic expression, but you can obtain the y-values directly from the graph for each x=-1 and x=1. What are they? f(1)=? f(-1)=?

OpenStudy (anonymous):

(-1,0) (1,0) ??

OpenStudy (mathmale):

Hold a moment, please.

OpenStudy (anonymous):

okay

OpenStudy (mathmale):

Jon: I did some quick calculations and have found that the equation of this parabola is y=(x-2)-3. Please evaluate this function at x=1 and again at x=-1.

OpenStudy (anonymous):

what do you mean evaluate it?

OpenStudy (mathmale):

"evaluate" is a command that means "find the value of." The equation of the parabola shown in your diagram is y=(x-2)^2-3. I am asking you to substitute x=1 and then x=-1 to determine the y value for each of those x values. That's what I mean when I ask you to "evaluate the function at 1 and -1."

OpenStudy (anonymous):

okay, give me a minute to work it out.

OpenStudy (anonymous):

y=6 and y=-2 ??

OpenStudy (mathmale):

Working on it; I need to check something; I'll be right with you.

OpenStudy (anonymous):

okay.

OpenStudy (mathmale):

Many apologies. My own equation for the parabola was wrong. It should be y=(x-1)^2-3. So sorry. Please evaluate this new equation at both 1 and -1.

OpenStudy (anonymous):

okay, one second....

OpenStudy (anonymous):

I got y=1 and y=-3

OpenStudy (mathmale):

Great. So, if x=1, y=-3, and if x=-1, y=1. Now please go back to that graph that you posted and determine for yourself whether or not the graph says the same thing about the values of y at x=1 and x=-1.

OpenStudy (mathmale):

Jon, sorry to have to squash that, but your initial results were correct and your later results were not. Please look at the graph as I've asked.

OpenStudy (mathmale):

Jon: Looking at the graph: If x = -1, what y value corresponds to that x?

OpenStudy (anonymous):

2 ??

OpenStudy (mathmale):

Locate x=-1 on the x-axis and then move straight up until you hit the graph. y=1 there. Please verify this for yourself. Now do the same for x=1. If x=1, y=??

OpenStudy (anonymous):

I'm so dumb. I forgot to fix the parabola equation on the graphing program I'm using....

OpenStudy (mathmale):

The formula for "average value of a function on the interval [a,b]" is\[Ave=\frac{ f(b)-f(a) }{ b-a }.\] In this case your b = 1 and your a = -1. Thus, you want \[Ave=\frac{ f(1)-f(-1) }{ 1-[-1] }\] Using the y -values you've just found algebraically and from the graph, please find Ave. Hint: f(1)=-3.

OpenStudy (mathmale):

Jon, you provided a nice illustration of this function when you presented the problem. If you're more comfortable using your calculator, fine, but I'd prefer you get the info (y-values) directly from the graph you initially gave me.

OpenStudy (mathmale):

Jon, hate to rush you, but we need to move faster on this. I need to get off the computer very soon. Please, check the graph again and see whether or not you agree on the following: y=1 when x=-1; y=-3 when x=1.

OpenStudy (mathmale):

Sorry, Jon, I'm getting off the computer now. If you're interested, try to complete this problem, and, if you wish, message me when you're ready to resume work. Nice working with you. So long.

OpenStudy (anonymous):

Thank you for helping!!!!! I really appreciate it.